PHOTOIONIZATION OF LIQUID HE DROPLETS

\underline{DARCY S. PETERKA}, LIONEL POISSON, College of Chemistry, UC Berkeley, Berkeley Ca, 94720; MUSAHID AHMED, Chemical Science Division, LBNL, Berkeley Ca, 94720; and DANIEL NEUMARK, College of Chemistry, UC Berkeley, Berkeley Ca, 94720.

We have studied the VUV photoionization process in helium droplets with Velocity Map Imaging at the Advanced Light Source. This multiplexed technique yields the velocity of the photoelectron, which can be used to understand the dynamics of the ionization process. Results will be presented for ionization processes below the energy of the He atom threshold.